OpenStudy (anonymous):

Solve: z2 + 2iz - 1 = 0

OpenStudy (anonymous):

-1=i^2 z^2+2iz+i^2=0 (z+i)^2=0 z=-i

OpenStudy (anonymous):

see thats what i got but i thought i couldnt be negative :/

OpenStudy (anonymous):

see thats what i got but i thought 'i' couldnt be negative

OpenStudy (anonymous):

-i perfectly exists.

OpenStudy (anonymous):

thanks!! i was about to flip wheni kept reworking it and still got -i